Target Costing

Laser Cast, Inc., manufactures color laser printers. Model J20 presently sells for $475 and has a total product cost of $380, as follows:

Direct materials $270
Direct labor 70
Factory overhead 40
Total $380

It is estimated that the competitive selling price for color laser printers of this type will drop to $450 next year. Laser Cast has established a target cost to maintain its historical markup percentage on product cost. Engineers have provided the following cost-reduction ideas:

  1. Purchase a plastic printer cover with snap-on assembly, rather than with screws. This will reduce the amount of direct labor by 9 minutes per unit.
  2. Add an inspection step that will add six minutes per unit of direct labor but reduce the materials cost by $9 per unit.
  3. Decrease the cycle time of the injection molding machine from four minutes to three minutes per part. Thirty percent of the direct labor and 42% of the factory overhead are related to running injection molding machines.

The direct labor rate is $29 per hour.

a. Determine the target cost for Model J20 assuming that the historical markup on product cost and selling price are maintained. Round your final answer to two decimal places.
$

b. Determine the required cost reduction. Enter as a positive number. Round your final answer to two decimal places.
$

c. Evaluate the three engineering improvements together to determine if the required cost reduction (drift) can be achieved. Enter all amounts as positive numbers. Do not round interim calculations but round your final answers to two decimal places.

1. Direct labor reduction $
2. Additional inspection $
3. Injection molding productivity improvement $
Total savings $
Answer #1

Solution:

a)

markup on Product cost = (sell price-cost)/cost

(475-380)/380

25.00%

Hence Target cost = sell price - markup

Hence Target cost = sell price -(25%*target cost)

let the target cost be x

x+0.25x=450

x=450/1.25

x=

360.00

b) required cost reduction = 380-360

   20.00

c)

calculation

$

1.) direct labor reduction

29*9/60

4.35

2.) Additional inspection

9 - (29*6/60)

6.01

3.) Injection molding productivity improvement

((70*0.30)+(40*0.42))*1/4

9.45

Total savings

19.81